1) Use the Acid-Base table to determine the pKa of the weak acid HCN.

2) What is the pH of a 0.00412 M solution of HCl?

3) What is the pH of 0.00685 M solution of NaOH?
Answer #1

ans 1

the weak acid HCN, Ka= 6.0x10-10,

pKa = -log Ka = 9.22.

ans 2

pH = -log [H+] = -log (0.00412)

pH=2.39

ans 3

pOH = -log [OH-] = -log (0.00685)

pOH=2.16

pH + pOH = 14

pH = 14-pOH = 14-2.16

pH = 11.8